The last dividend paid by Coppard Inc. was $1.25. The dividend growth rate is expected to be constant at 42.5% for 3 years, after which dividends are expected to grow at a rate of 6% forever. If the firm's required return (rs) is 11%, what is its current stock price?
Answer #1
As per dividend discount method, current stock price is the present value of dividends.
Step-1:Present value of next 3 year's dividend
Year Dividend Present value of 1 Present value of dividend
a b c=1.11^-a d=b*c
1 $       1.78                 0.9009 $                  1.60
2 $       2.54                 0.8116 $                  2.06
3 $       3.62                 0.7312 $                  2.65
Total $                  6.31
Working:
Year 1 dividend = 1.25*1.425 = $       1.78
Year 2 dividend = 1.78*1.425 = $       2.54
Year 3 dividend = 2.54*1.425 = $       3.62
Step-2:Terminal value of dividend at the end of year 3
Terminal Value = D3*(1+g)/(Ke-g) Where,
= 3.62*(1+0.06)/(0.11-0.06) D3 Year 3 dividend $       3.62
= $                76.74 g Growth rate 6%
Ke Required Return 11%
Step-3:Present value of terminal value
Present value = Terminal value at the end of Year 3*Present value of 1
= $                76.74 *      0.7312
= $                56.11
Working:
Present value of 1 = (1+i)^-n Where,
= (1+0.11)^-3 i 11%
=                  0.7312 n 3
Step-4:Present value of all dividends
Present value of all dividends = $       6.31 + $    56.11
= $    62.42
As present value of dividends is $ 62.42. So, current stock price is $ 62.42.